HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SPECTROSCOPIC STUDY OF BUFFER GAS COOLED METHANOL IN THE 2OH STRETCHING RANGE

\underline{A. LIBERT} , Institute of Condensed Matter and Nanosciences, Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium AND Royal Belgian Institute for Space Aeronomy, Brussels, Belgium; A. ROUCOU, Université du

Littoral Côte d’Opale, UR4493, LPCA, Laboratoire de Physico-Chimie de l’Atmosphère, F-59140 Dunkerque, France; B. HAYS, R. GLORIEUX, Institute of Condensed Matter and Nanosciences, Université catholique de Louvain, B-1348

Louvain-la-Neuve, Belgium; S. ROBERT, Royal Belgian Institute for Space Aeronomy, Brussels, Belgium; B. FABRE, Université de Bordeaux, CNRS, CEA, CELIA, UMR5107, F33405 Talence, France; S. KASSI, Univ. Grenoble Alpes,

CNRS, LIPhy, Grenoble, France; X. URBAIN, C. LAUZIN, Institute of Condensed

Matter and Nanosciences, Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

Methanol (CH_{3}OH) is a six-atoms asymmetric top molecule presenting internal rotation. The size of this molecule and the presence of a large amplitude motion lead to a dense rotational structure. This structure gets even more complicated when one goes up in energy with vibrational excitations. Due to its complicated spectrum, this molecule remains poorly known, especially in the near-infrared. This frequency range was recently explored by Svoboda et al.(Phys. Chem. Chem. Phys., 17 (2015) 15710) and by Rakvský et al. (Phys. Chem. Chem. Phys., 23 (2021) 20193), probing the 2OH stretching vibration overtone around 7200 cm^{−1}.

In this contributed talk, we will present the measurement of the full 2ν1 vibration overtone around 7200 cm^{−1} cooled down to 20 K with a buffer gas cooling setup coupled to a Cavity ring down spectrometer with a sensitivity of α_{min} = 2×10^{−10}cm^{−1}. The two elements of the built in-house apparatus are first briefly introduced. The assignment procedure is then detailed. It is based on the extension of the ground state combination differences method (GSCD) to all the allowed transitions that lead to a same upper energy state. Indeed, these transitions form a unique pattern spaced on the wavenumber axis by the differences in energy between the involved lower energy states. This method was first presented by Rakvský et al. We discuss the use of the intensities as an additional criterion to decide between correct assignments and fortuitous detection of the pattern on the experimental spectrum.

p-number: p205  Submitted on Thu, 15 Jun 2023 22:09:01 +0200